# Guiding-center dynamics in a screw-pinch magnetic field

### v1

@@ III.2 Newton’s equations of motion in screw-pinch magnetic geometry
Equation (28), therefore, yields the three coupled equations for the velocity components $(v_{\|},v_{\bot},\zeta)$:
@@
where the Frenet-Serret coefficients $(\kappa,\tau,\tau_{m})$ are given by Eqs. (14) and (16) and the gyrogauge component $R_{\bot}\equiv 1/r-\kappa$ is given by Eq. (17). Using Eqs. (24) and (34), we note that, while the time derivative of the lowest-order magnetic moment $\mu_{0}\equiv\frac{1}{2}\epsilon^{2}v_{\bot}^{2}\cos\Theta$:
$\frac{d\mu_{0}}{dt}$ $=$ $\epsilon^{2}v_{\bot}\,\cos\Theta\;\frac{dv_{\bot}}{dt}\;-\;\frac{\epsilon^{2}}{2}\,v_{\bot}^{2}\,\sin\Theta\;\frac{d\Theta}{dt}$ (36)
$=$ $\epsilon^{3}\,v_{\|}\,v_{\bot}\,\cos\Theta\;\sin\zeta\left[v_{\|}\,\kappa\;+\;v_{\bot}\,(\tau-\Theta^{\prime})\,\cos\zeta\right]\;-\;\frac{\epsilon^{3}}{2}\,v_{\bot}^{3}\,\Theta^{\prime}\,\sin\Theta\;\sin\zeta,$
@@
Lastly, we note that the kinetic energy $E=\epsilon^{2}(v_{\|}^{2}+v_{\bot}^{2})/2$ is a constant of motion, since $dE/dt=\epsilon\,v_{\|}\,dv_{\|}/dt+\epsilon\,v_{\bot}\,dv_{\bot}/dt=0$. The parallel velocity $v_{\|}=v\,\cos\lambda$ and the perpendicular velocity $v_{\bot}=v\,\sin\lambda$ can therefore be expressed in terms of the pitch angle $\lambda$, so that the equation of motion for the pitch angle $\lambda$ is expressed as
@@ IV Particle Lagrangian Mechanics
As an alternative to the geometric Newtonian representation presented in Sec. III, we now explore the Lagrangian representation of the particle dynamics in the screw-pinch magnetic field (8). We now choose the magnitude of the magnetic field to be a function of $\Theta$ only (i.e., it is independent of $z$):
$B(\Theta)\;\equiv\;\sec\Theta,$ (38)
where we note that $\nabla\ln B=-\,\Theta^{\prime}\,\tan\,\Theta\;\widehat{\sf e}_{1}$. We note that, while the choice (38) allows us to obtain explicit expressions, it does not impact on the main results of our work.
@@ IV.2 Particle Lagrangian
Using the normalized magnetic vector potential ${\bf A}=\psi\;\nabla\theta-\Psi\;\nabla z$ associated with the magnetic field (39), we express the normalized particle Lagrangian in cylindrical coordinates as
@@
which follows after substituting Eq. (34) for $dv_{\bot}/dt$. Hence, as expected, the Euler-Lagrange equations exactly agree with Newton’s equations.
@@ IV.3 Reduced radial particle dynamics
$V(r)\;=\;\frac{1}{2}\,\left[\left(\Psi-\overline{\Psi}\right)^{2}\;+\;\frac{1}{r^{2}}\,\left(\psi-\overline{\psi}\right)^{2}\right].$ (51)
@@
The reduced radial action integral
$J_{r}(E,P_{\theta},P_{z})\;\equiv\;\frac{1}{2\pi}\oint P_{r}\,dr\;=\;\frac{1}{2\pi}\oint\sqrt{2\left[E\;-\;V(r)\right]}\;dr$ (53)
is also an invariant of the reduced radial motion. With three separate exact constants of motion $(E,P_{\theta},P_{z})$, the motion of a charged particle in this screw-pinch magnetic field is said to be integrable. We will return to the reduced radial action integral (53) in Sec. VI to show that it matches exactly with the magnetic moment up to third order in $\epsilon$.
@@ V.1 Zeroth-order magnetic moment
The lowest-order magnetic moment $\mu_{0}=m\,|d{\bf r}_{\bot}/dt|^{2}/2\Omega$ (written in units of action) is defined as
@@ V.2 First-order magnetic moment
$\mu_{1}\;=\;\left(\overline{\mu_{0}}\,\nabla\ln B+\epsilon^{2}\,v_{\|0}^{2}\,\cos\Theta\;\mbox{$\kappa$}\right)\,\mbox{$\cdot$}\,\mbox{$\rho$}_{0}\;-\;\overline{\mu_{0}}\,v_{\|0}\,\cos\Theta\;(\tau_{m}+\alpha_{1}),$ (56)
@@
$\mbox{$\rho$}_{0}$ $=$ $\frac{\widehat{{\sf b}}}{B}\,\mbox{$\times$}\,{\bf v}_{\bot}\;=\;v_{\bot 0}\,\cos\Theta_{0}\left(\cos\zeta\,\widehat{\sf e}_{1}\;-\;\sin\zeta\,\widehat{\sf e}_{2}\right),$ (57)
and the torsional correction
$\tau_{m}+\alpha_{1}\;\equiv\;\frac{3}{2}\tau_{m}-\widehat{\bot}\,\mbox{$\cdot$}\,\nabla\widehat{{\sf b}}\,\mbox{$\cdot$}\,\widehat{\rho}\;=\;\tau\left(\frac{3}{2}-\cos^{2}\zeta\right)\;+\;\Theta^{\prime}\left(\frac{3}{2}-\sin^{2}\zeta\right)$ (58)
is expressed in terms of the Frenet-Serret torsion $\tau$ and the magnetic-pitch shear $\Theta^{\prime}$. Equation (56) may be rewritten as
$\mu_{1}\;=\;\overline{\mu_{0}}\,\mbox{$\rho$}_{0}\,\mbox{$\cdot$}\,\nabla\ln B\;+\;\rho_{\|0}\left[\epsilon^{2}\,v_{\|0}\,\mbox{$\rho$}_{0}\,\mbox{$\cdot$}\,\mbox{$\kappa$}\;-\;\overline{\mu_{0}}\;(\tau_{m}+\alpha_{1})\right],$ (59)
where the first term is
$\overline{\mu_{0}}\,\mbox{$\rho$}_{0}\,\mbox{$\cdot$}\,\nabla\ln B\;=\;-\;\overline{\mu_{0}}\,\rho_{\bot 0}\cos\zeta\;\left(\Theta^{\prime}_{0}\;\tan\Theta_{0}\right).$
We now use the inverse guiding-center transformation $z^{\alpha}\equiv Z_{\rm gc}^{\alpha}-\epsilon\,G_{1}^{\alpha}+\cdots$ presented in App. A, so that we find
@@
Hence, the full magnetic moment is
@@
which implies that the first-order corrections of $\mu_{0}$ are exactly canceled out by $\mu_{1}$.
@@ VI Expansion of the Reduced Radial Action
$J_{r}\;=\;\frac{1}{2\pi}\oint P_{r}\,dr\;=\;\frac{1}{2\pi}\int_{0}^{2\pi}\frac{dr}{dt}\;\frac{\partial r}{\partial\zeta}\;d\zeta\;=\;\frac{1}{2\pi}\int_{0}^{2\pi}\left(\frac{dr}{dt}\right)^{2}\;\frac{d\zeta}{(d\zeta/dt)},$ (66)
where $(dr/dt,d\zeta/dt)$ are given by Eqs. (24) and (35), respectively. In order to compare the reduced radial action integral (66) with the magnetic moment (65), we need to expand $(dr/dt)^{2}$ up to third order in $\epsilon$ and $(d\zeta/dt)^{-1}$ up to first order in $\epsilon$. First, we find
@@
where the last term $(\cdots)$ combines contributions that are proportional to $\sin^{2}\zeta\,\cos\zeta$.
@@
we obtain the final result
@@
which confirms that, up to third order in $\epsilon$, the reduced radial action integral is identical to the magnetic moment.
@@ VII Conclusion and Discussion
By using the geometric formulation of the Newtonian mechanics of charged particle motion in a screw-pinch magnetic field, we have demonstrated Kruskal’s identity between the reduced radial action integral (66), from which the Kruskal adiabatic-invariant series is derived, and the magnetic moment (65), which is expanded explicitly in powers of the magnetic-field non-uniformity. In contrast to the previous work by Holas et al.Holas et al. (2025), we have explicitly demonstrated the Kruskal identity up to third order in $\epsilon$. By exploiting the algorithmic nature of the Lie-transform approach to guiding-center dynamics, which would generate explicit expressions for $\mu_{2}$ and $(r_{2},\lambda_{2})$, it is possible to prove the Kruskal identity $J_{r}=\mu_{\rm gc}$ up to fourth order in $\epsilon$ (i.e., second order in magnetic-field non-uniformity).
@@
$\mu_{\rm gy}\;\equiv\;\frac{\epsilon}{2\pi}\int_{0}^{2\pi}{\bf P}_{\rm gy}\,\mbox{$\cdot$}\,\frac{\partial\mbox{$\rho$}_{\rm gy}}{\partial\zeta}\;d\zeta.$ (72)
where an additional time-dependent transformation
@@
where $\delta$ is defined in terms of the amplitude of the field perturbationBrizard and Hahm (2007) and the Lie-transform operator ${\sf T}_{n}^{-1}\equiv\exp(\delta^{n}{\mathcal{L}}_{n})$ is defined in terms of the $n$th-order Lie derivative ${\mathcal{L}}_{n}\equiv{\sf G}_{n}\cdot{\sf d}$ generated by the vector field ${\sf G}_{n}$, whose purpose is the remove the gyroangle dependence from the gyrocenter dynamics (i.e., the perturbed guiding-center dynamics).
In future work, we could explore whether the bounce-center dynamics, with the bounce-center action

### v2

@@ III.2 Newton’s equations of motion in screw-pinch magnetic geometry
Equation (28), therefore, yields the three coupled equations for the velocity coordinates:
@@
where the Frenet-Serret coefficients $(\kappa,\tau,\tau_{m})$ are given by Eqs. (14) and (16) and the gyrogauge component $R_{\bot}\equiv 1/r-\kappa$ is given by Eq. (17). Using Eqs. (24) and (34), we note that, while the time derivative of the lowest-order magnetic moment $\mu_{0}\equiv\frac{1}{2}\epsilon^{2}v_{\bot}^{2}/B(\Theta)$:
$\frac{d\mu_{0}}{dt}$ $=$ $\frac{\epsilon^{2}v_{\bot}}{B(\Theta)}\;\frac{dv_{\bot}}{dt}\;-\;\frac{\epsilon^{2}\,v_{\bot}^{2}\,B^{\prime}(\Theta)}{2\,B^{2}(\Theta)}\;\frac{d\Theta}{dt}$ (36)
$=$ $\frac{\epsilon^{3}\,v_{\|}\,v_{\bot}}{B(\Theta)}\;\sin\zeta\left[v_{\|}\,\kappa\;+\;v_{\bot}\,(\tau-\Theta^{\prime})\,\cos\zeta\right]\;-\;\frac{\epsilon^{3}\,v_{\bot}^{3}\,B^{\prime}(\Theta)}{2\,B^{2}(\Theta)}\;\Theta^{\prime}\;\sin\zeta,$
@@
Lastly, we note that the kinetic energy $E=\epsilon^{2}(v_{\|}^{2}+v_{\bot}^{2})/2$ is a constant of motion, since $dE/dt=\epsilon^{2}\,v_{\|}\,dv_{\|}/dt+\epsilon^{2}\,v_{\bot}\,dv_{\bot}/dt=0$ when Eqs. (33)-(34) are taken into account. The parallel velocity $v_{\|}=v\,\cos\lambda$ and the perpendicular velocity $v_{\bot}=v\,\sin\lambda$ can, therefore, be expressed in terms of the pitch angle $\lambda$, so that the equation of motion for the pitch angle $\lambda$ is expressed as
@@ IV Particle Lagrangian Mechanics
As an alternative to the geometric Newtonian representation presented in Sec. III, we now explore the Lagrangian representation of the particle dynamics in the screw-pinch magnetic field (8). We now choose the magnitude of the magnetic field to be defined as
$B(\Theta)\;\equiv\;\sec\Theta.$ (38)
We note that, while the choice (38) allows us to obtain explicit expressions, it does not impact on the main results of our work. While the Lagrangian formalism yields identical results obtained from the Newtonian formalism adopted in Sec. III, it is presented, here, as a way to derive the reduced radial action integral [see Eq. (53)], which can then be inserted in the Kruskal identity (6) for the screw-pinch magnetic geometry.
@@ IV.2 Particle Lagrangian
Using the normalized magnetic vector potential ${\bf A}=\psi\;\nabla\theta-\Psi\;\nabla z$ associated with the magnetic field (39), we express the particle Lagrangian in cylindrical coordinates as
@@
which follows after substituting Eq. (34) for $dv_{\bot}/dt$. Hence, as expected, the Euler-Lagrange equations exactly agree with Newton’s equations. While this result appears trivial, it serves as an explicit proof of the validity of the geometric Newtonian formalism introduced in Sec. III.
@@ IV.3 Reduced radial particle dynamics
$V(r)\;=\;\frac{1}{2}\,\left[\left(\Psi(\psi)-\overline{\Psi}\right)^{2}\;+\;\frac{1}{r^{2}}\,\left(\psi(r)-\overline{\psi}\right)^{2}\right].$ (51)
@@
The reduced radial action integralGoldstein et al. (2002)
$J_{r}(E,P_{\theta},P_{z})\;\equiv\;\frac{1}{2\pi}\oint P_{r}\,dr\;=\;\frac{1}{2\pi}\int_{0}^{2\pi}\sqrt{2\left[E\;-\;V(r)\right]}\;\frac{dr}{d\zeta}\;d\zeta$ (53)
is also an invariant of the reduced radial motion. With three separate exact constants of motion $(E,P_{\theta},P_{z})$, the motion of a charged particle in this screw-pinch magnetic field is said to be integrable. In Eq. (53), using the equivalence of the Newtonian and Lagrangian formalisms, we will express $dr/d\zeta\equiv(dr/dt)/(d\zeta/dt)$ in terms of $dr/dt=\sqrt{2\,[E-V(r)]}$ and $d\zeta/dt$, which are defined respectively by the Newtonian expressions (24) and (35) . The benefit of the Lagrangian formalism is that it provides an explicit derivation of the reduced radial potential (51), which yields the reduced radial action integral (53). However, because the azimuthal magnetic flux $\Psi(\psi)$ cannot be solved in closed form, we will show that the geometric Newtonian formalism will yield explicit expressions for the reduced radial action integral (53). We will return to the reduced radial action integral (53) in Sec. VI to show that it matches exactly with the magnetic moment up to third order in $\epsilon$.
@@ V.1 Zeroth-order magnetic moment
The lowest-order (normalized) magnetic moment $\mu_{0}=|d{\bf r}_{\bot}/dt|^{2}/2B$ is defined as
@@ V.2 First-order magnetic moment
$\mu_{1}\;=\;\left(\overline{\mu_{0}}\,\nabla\ln B+\epsilon^{2}\,v_{\|0}^{2}\,\cos\Theta\;\mbox{$\kappa$}\right)\,\mbox{$\cdot$}\,\mbox{$\rho$}_{0}\;-\;\epsilon\,\overline{\mu_{0}}\,v_{\|0}\,\cos\Theta\;(\tau_{m}+\alpha_{1}),$ (56)
@@
$\mbox{$\rho$}_{0}$ $=$ $\frac{\widehat{{\sf b}}}{B}\,\mbox{$\times$}\,\,\epsilon\;{\bf v}_{\bot}\;=\;\epsilon\,v_{\bot 0}\,\cos\Theta_{0}\left(\cos\zeta\,\widehat{\sf e}_{1}\;-\;\sin\zeta\,\widehat{\sf e}_{2}\right),$ (57)
[…]
$\alpha_{1}\;=\;\frac{1}{2}\,(\Theta^{\prime}-\tau)\,\cos(2\zeta)\;=\;\frac{1}{2}\,\tau_{m}\;-\;\left(\tau\;\cos^{2}\zeta\;+\;\Theta^{\prime}\;\sin^{2}\zeta\right),$
so that the torsional correction
$\tau_{m}+\alpha_{1}\;=\;\tau_{0}\left(\frac{3}{2}-\cos^{2}\zeta\right)\;+\;\Theta^{\prime}_{0}\left(\frac{3}{2}-\sin^{2}\zeta\right)$ (58)
is expressed in terms of the Frenet-Serret torsion $\tau$ and the magnetic-pitch shear $\Theta^{\prime}$, which are both evaluated at $r_{0}$. Equation (56) may, therefore, be rewritten as
$\mu_{1}\;=\;\overline{\mu_{0}}\,\mbox{$\rho$}_{0}\,\mbox{$\cdot$}\,\nabla\ln B\;+\;\epsilon\,\rho_{\|0}\left[\epsilon^{2}\,v_{\|0}\,\mbox{$\rho$}_{0}\,\mbox{$\cdot$}\,\mbox{$\kappa$}\;-\;\overline{\mu_{0}}\;(\tau_{m}+\alpha_{1})\right],$ (59)
where the first term is $\overline{\mu_{0}}\,\mbox{$\rho$}_{0}\,\mbox{$\cdot$}\,\nabla\ln B=-\;\epsilon\,\overline{\mu_{0}}\,\rho_{\bot 0}\cos\zeta\;\left(\Theta^{\prime}_{0}\;\tan\Theta_{0}\right)$.
We now use the inverse guiding-center transformation $z^{\alpha}\equiv Z_{\rm gc}^{\alpha}-\epsilon\,G_{1}^{\alpha}+\cdots$ presented in App. B, which yields
@@
Hence, by combining Eqs. (55) and (64), we obtain the full magnetic moment (up to third order in $\epsilon$)
@@
which implies that the first-order corrections of $\mu_{0}$ are exactly canceled out by $\mu_{1}$. We note that this cancellation is not unique to the case of the screw-pinch magnetic geometry, as it is also observed in the case of a slab magnetic fieldHolas et al. (2025) (see also App. A in this paper).
@@ VI Expansion of the Reduced Radial Action
$J_{r}\;=\;\frac{1}{2\pi}\oint P_{r}\,dr\;=\;\frac{1}{2\pi}\int_{0}^{2\pi}\left(\frac{dr}{dt}\right)^{2}\;\frac{d\zeta}{(d\zeta/dt)},$ (66)
where the Newtonian equations for $(dr/dt,d\zeta/dt)$ are given by Eqs. (24) and (35), respectively. In order to compare the reduced radial action integral (66) with the magnetic moment (65), we need to expand $(dr/dt)^{2}$ up to third order in $\epsilon$ and $(d\zeta/dt)^{-1}$ up to first order in $\epsilon$. First, we find
@@
where gyroangle averaging is denoted as $\langle\cdots\rangle$ and the last term $(\cdots)$ in Eq. (70) combines contributions that are proportional to $\sin^{2}\zeta\,\cos\zeta$.
@@
we obtain the final result from Eq. (70):
@@
which confirms that, up to third order in $\epsilon$, the reduced radial action integral is identical to the magnetic moment (65).
@@ VII Conclusion and Discussion
The present work was concerned with an explicit proof of the Kruskal identity (6) in a doubly-symmetric screw-pinch magnetic geometry. In recent work by Holas et al.Holas et al. (2025), a Lagrangian formulation was used to derive an explicit expression for the reduced radial action integral (66), which is expressed in terms of the axial magnetic flux $\psi(r)$ and the azimuthal magnetic flux $\Psi(\psi)$ defined in Eq. (40). Unfortunately, since the azimuthal magnetic flux $\Psi(\psi)$ cannot be solved in closed form, the Lagrangian formalism adopted by Holas et al.Holas et al. (2025) yielded results that could only prove the Kruskal identity through computer algebra up to third order in $\epsilon$. By using the geometric formulation of the Newtonian mechanics of charged particle motion in a screw-pinch magnetic field in the present work, we have demonstrated Kruskal’s identity between the reduced radial action integral (66), from which the Kruskal adiabatic-invariant series is derived, and the magnetic moment (65), which is expanded explicitly in powers of the magnetic-field non-uniformity. By exploiting the algorithmic nature of the Lie-transform approach to guiding-center dynamics, which would generate explicit expressions for $\mu_{2}$ and $(r_{2},\lambda_{2})$, it is possible to prove the Kruskal identity $J_{r}=\mu_{\rm gc}$ up to fourth order in $\epsilon$ (i.e., second order in magnetic-field non-uniformity).
@@
$\mu_{\rm gy}\;\equiv\;\frac{\epsilon}{2\pi}\int_{0}^{2\pi}{\bf P}_{\rm gy}\,\mbox{$\cdot$}\,\frac{\partial\mbox{$\rho$}_{\rm gy}}{\partial\zeta}\;d\zeta\;=\;\mu_{\rm gc}\;+\;\delta\,\mu_{\rm gy1}\;+\;\cdots,$ (72)
where $\mu_{\rm gc}$ denotes the possibly non-perturbative (unperturbed) guiding-center magnetic moment, the ordering parameter $\delta$ is defined in terms of the amplitude of the field perturbationBrizard and Hahm (2007), and an additional time-dependent transformation
@@
is defined in terms of the Lie-transform operator ${\sf T}_{n}^{-1}\equiv\exp(-\delta^{n}{\mathcal{L}}_{n})$, with the $n$th-order Lie derivative ${\mathcal{L}}_{n}\equiv{\sf G}_{n}\cdot{\sf d}$ generated by the vector field ${\sf G}_{n}$, whose purpose is the remove the gyroangle dependence from the gyrocenter dynamics (i.e., the perturbed guiding-center dynamics). It is not entirely clear whether the existence of a non-perturbative integral expression for the gyrocenter magnetic moment (72) has any relevance for practical gyrokinetic numerical simulations. However, the statistical full-orbit analysis used by Burby et al.J. W. Burby et al. (2025) might still be able to yield some interesting results when applied to the perturbed guiding-center dynamics, which could be explored in future work.
In additional future work, we could explore whether the bounce-center dynamicsLittlejohn (1982b); Cary and Brizard (2009), with the bounce-center action defined as